OpenStudy (gamereternity):

A customer at an online movie rental company pays a total of $25 after 1 month, a total of $35 after 2 months, and a total of $45 after 3 months. The total amount paid is represented by T and the number of months by m. Which equation models the total payment, T, after m months? A. T = 10m + 15 B. T = 15m + 10 C. T = 25m D. T = 5m + 30

OpenStudy (mightychondrion):

Only model A fits the values given. 25=10*(1) + 15 35=10*(2) + 15 45=10*(3) +15

OpenStudy (gamereternity):

ty